\begin{thebibliography}{MMM}
\bibitem{AWR17} J. Altschuler, J.  Weed and P. Rigollet, Near-linear time approximation algorithms for optimal transport via Sinkhorn iteration,
NIPS'17: Proceedings of the 31st International Conference on Neural Information Processing Systems, December 2017,  1961-1971.

\bibitem{ACB17} M. Arjovsky, S. Chintala and L. Bottou, Wasserstein Generative Adversarial Networks, Proceedings of the 34th International Conference on Machine Learning, PMLR 70:214-223, 2017.
  
 \bibitem{BCN19} J-D. Benamou, G. Carlier, and L. Nenna. Generalized incompressible 
ows, multi-marginal transport and Sinkhorn algorithm. Numerische Mathematik, 142(1):(2019)33-54.
 
\bibitem{BGKL17} J. Bigot, R. Gouet, T. Klein, and A. L\'opez, Geodesic PCA in the Wasserstein space by convex PCA. Ann. Inst. H. Poincaré Probab. Statist., 53(1):1-26, 02 2017.

\bibitem{BPPH11} N. Bonneel, M. van de Panne, S. Paris, and W. Heidrich. Displacement interpolation using
Lagrangian mass transport. ACM Trans. Graph., 30(6):158:1-158:12, December 2011.
 
\bibitem{CEFZ23} S. Cole, M.  Eckstein, S. Friedland, K. {\.Z}yczkowski, On Quantum Optimal Transport, \emph{Mathematical Physics, Analysis and Geometry} (2023) 26:114,  67 pages,https://doi.org/10.1007/s11040-023-09456-7
 
 
\bibitem{Cut13}  M. Cuturi, Sinkhorn distances: Lightspeed computation of optimal transport, In C. J. C. Burges, L. Bottou, M. Welling, Z. Ghahramani, and K. Q. Weinberger, editors, Advances in Neural
Information Processing Systems 26, pages 2292-2300, Curran Associates, Inc., 2013..

\bibitem{FCCR18} R. Flamary, M. Cuturi, N. Courty, A. Rakotomamonjy, Wasserstein Discriminant Analysis, Machine Learning  107 (2018), 1923--1945.



\bibitem{Fri20} S. Friedland,  Tensor optimal transport, distance between sets of measures and tensor scaling, arXiv:2005.00945, version 1, 2020.

\bibitem{FECZ22}  S. Friedland, M. Eckstein, S. Cole, K. {\.Z}yczkowski, Quantum Monge-Kantorovich problem and transport distance between density matrices,  Physical Review Letters 129,  Issue 11, 110402 - Published 7 September 2022.

\bibitem{FV18} G. Friesecke,  and D.  V\"ogler,  Breaking the curse of dimension in multi-marginal Kantorovich optimal transport on finite state spaces, \emph{SIAM J. Math. Anal. } 50 (2018), no. 4, 3996-4019.

\bibitem{HRCK} I Haasler, A Ringh, Y Chen, J Karlsson,
Multimarginal optimal transport with a tree-structured cost and the Schr\"odinger bridge problem, {\it SIAM J. Control Optim. } 59 (2021), no. 4, 2428-2453.
 
\bibitem{Kan} L.V Kantorovich, On the translocation of masses. Dokl. Akad. Nauk.
USSR 37 (1942), 199–201. English translation in J. Math. Sci. 133, 4 (2006),
1381–1382.

\bibitem{LZZ17} Z. Li,  F. Zhang and X.-D.  Zhang,  On the number of vertices of the stochastic tensor polytope, {\it Linear Multilinear Algebra} 65 (2017), no. 10, 2064-2075.

\bibitem{LHCJ22}  
T. Lin, N. Ho,  M. Cuturi and M. I. Jordan, 
On the complexity of approximating multimarginal optimal transport
{\it J. Mach. Learn. Res. } 23 (2022), Paper No. [65], 43 pp.
 
 \bibitem{LG15} J. R. Lloyd and Z. Ghahramani, Statistical model criticism using kernel two sample tests, In Proceedings of the 28th International Conference on Neural Information Processing Systems, NIPS’15, pages 829–837, Cambridge, MA, USA, 2015. MIT Press.
 
 \bibitem{MJ15} J. Mueller and T. Jaakkola, Principal differences analysis: Interpretable characterization of differences between distributions, In Proceedings of the 28th International Conference on Neural Information Processing Systems, NIPS’15, pages 1702–1710, Cambridge, MA, USA, 2015. MIT Press.


\bibitem{NN94} Yu. Nesterov and A. S. Nemirovski, Interior-Point Polynomial Algorithms in Convex Programming, {\it Stud. Appl. Math.}, SIAM, Philadelphia, 1994.

 \bibitem{PZ16} V. M. Panaretos and Y. Zemel, Amplitude and phase variation of point processes, Ann. Statist., 44(2):771–812, 04 2016.
 
 
 \bibitem{PW09} O. Pele and M.Werman, Fast and robust earth mover’s distances, in 2009 IEEE 12th International Conference on Computer Vision, 460-467, Sept 2009.
 
 \bibitem{Pi68} W.P.  Pierskalla, The multidimensional assignment problem,  \emph{Oper. Res.} 16 (1968),   422--431.
 
\bibitem{Po94}  A.B. Poore, Aubrey, Multidimensional assignment formulation of data association problems arising from multitarget and multisensor tracking, \emph{Comput. Optim. Appl. } 3 (1994), no. 1,  27-57. 
 
 \bibitem{Ren01} J. Renegar, {\it A Mathematical View of Interior-Point Methods}, in Convex Optimization, MOS-SIAM Ser. Optim., SIAM, Philadelphia, 2001.

 
\bibitem{RTG00} Y. Rubner, C. Tomasi, and L. J. Guibas, The earth mover’s distance as a metric for image retrieval,
\emph{Int. J. Comput. Vision}, 40(2):99–121, November 2000.

\bibitem{SL11} R. Sandler and M. Lindenbau, Nonnegative matrix factorization with earth mover’s distance metric for image analysis, IEEE Transactions on Pattern Analysis and Machine Intelligence, 33(8):1590–1602, Aug 2011.


\bibitem{SR04} G. J. Sz\'ekely and M. L. Rizzo. Testing for equal distributions in high dimension. Inter-Stat (London), 11(5):1-16, 2004.

\bibitem{TDGU20} N. Tupitsa, P. Dvurechensky, A. Gasnikov, and C. A. Uribe, Multimarginal optimal transport by accelerated alternating minimization,  59th IEEE Conference on Decision and Control, 6132-6137,IEEE, 2020.

\bibitem{VB96} L. Vandenberghe and S. Boyd, Semidefinite Programming, SIAM Review 38, March 1996, pp. 49-95.

\bibitem{Vil03}  C. Villani, Topics in optimal transportation, Graduate Studies in Mathematics, 58. American Mathematical Society, Providence, RI, 2003. .

\bibitem{Vil09} C. Villani, Optimal transport, Old and new, Grundlehren der Mathematischen Wissenschaften, 338, Springer-Verlag, Berlin, 2009.

\bibitem{YFFKNN} M. Yamashita, K. Fujisawa, M. Fukuda, K. Kobayashi, K. Nakata, and M. Nakata,{\it Latest developments in the SDPA family for solving large-scale SDPs}, in Handbook on Semidefinite, Conic and Polynomial Optimization, M. F. Anjos and J. B. Lasserre, eds.,Springer, New York, 2012, pp. 687-713.

 \end{thebibliography}

\documentclass[11pt]{amsart}

\usepackage{hyperref}

\usepackage{graphicx, enumerate, url}
\usepackage{amssymb,mathtools}
\usepackage{algorithm}
\usepackage{algpseudocode}
\usepackage{color}
\usepackage{tikz}
\numberwithin{equation}{section}
\numberwithin{algorithm}{section}

\theoremstyle{plain}
\newtheorem{theorem}{Theorem}[section]

\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{corollary}[theorem]{Corollary}

\newtheorem{definition}[theorem]{Definition}

\theoremstyle{remark}

\DeclareMathOperator{\tr}{Tr}
\DeclareMathOperator{\diag}{diag}

\newcommand\defeq{\stackrel{\mathclap{\normalfont\mbox{def}}}{=}}
\newcommand{\N}{\mathbb{N}}

\newcommand{\R}{\mathbb{R}}

\newcommand{\x}{\mathbf{x}}

\newcommand{\y}{\mathbf{y}}

\newcommand{\U}{\mathbf{U}}

\newcommand{\cC}{\mathcal{C}}

\newcommand{\cJ}{\mathcal{J}}

\newcommand{\cU}{\mathcal{U}}
\newcommand{\cV}{\mathcal{V}}
\newcommand{\cW}{\mathcal{W}}
\newcommand{\cX}{\mathcal{X}}

\newcommand{\be}{\mathbf{e}}
\newcommand{\bbf}{\mathbf{f}}
\newcommand{\bg}{\mathbf{g}}
\newcommand{\bh}{\mathbf{h}}
\newcommand{\bp}{\mathbf{p}}
\newcommand{\bq}{\mathbf{q}}

\newcommand{\bu}{\mathbf{u}}
\newcommand{\bv}{\mathbf{v}}
\newcommand{\bw}{\mathbf{w}}
\newcommand{\bx}{\mathbf{x}}
\newcommand{\by}{\mathbf{y}}

\newcommand{\rB}{\mathrm{B}}
\newcommand{\rC}{\mathrm{C}}
\newcommand{\rD}{\mathrm{D}}

\newcommand{\rL}{\mathrm{L}}
\newcommand{\rS}{\mathrm{S}}
\newcommand{\rU}{\mathrm{U}}
\newcommand{\rV}{\mathrm{V}}

\newcommand{\0}{\mathbf{0}}
\newcommand{\1}{\mathbf{1}}

\begin{document}
\title[IPM in tensor optimal transport]{Interior point method in tensor optimal transport}
\author[Shmuel Friedland]{Shmuel~Friedland}
\address{Department of Mathematics and Computer Science, University of Illinois at Chicago, Chicago, Illinois, 60607-7045, USA }
\email{friedlan@uic.edu}
\date{October 29, 2023}
\begin{abstract}  
We study a tensor optimal transport (TOT) problem for $d\ge 2$ discrete  measures. This is a linear programming problem on $d$-tensors.     We introduces an interior point method (ipm) for $d$-TOT with a corresponding barrier function.   Using  a "short-step" ipm following central path within $\varepsilon$ precision we estimate the number of iterations. 
\end{abstract}

\keywords{Tensor optimal transport, interior point method}

\subjclass[2010]{15A69, 52A41, 62H17, 65D19,  65K05, 90C25}
\maketitle
\section{Introduction} \label{sec:intro}
For $i\in\{1,2\}$ let $X_i$ be a random variables on $\Omega_i$ which has a finite number of values: $X_i:\Omega_i\to [n_i]$, where $[n]\defeq\{1,\ldots,n\}\subset \mathbb{N}$.  
 Assume that $\bp_i=(p_{1,i},\ldots,p_{n_i,i})$ is the column probability vector that gives the distribution of $X_i$: $\mathbb{P}(X_i=j)=p_{j,i}$.
Then the discrete Kantorovich optimal transport problem (OT) can be states as follows \cite{Kan}.  (See \cite{Vil03,Vil09} for modern account of OT.)  Let $Z$ be a random variable $Z: \Omega_1\times \Omega_2\to [n_1]\times[n_2]$ with contingency matrix (table)  $U\in\R_+^{n_1\times n_2}$ that gives the distribution of $Z$:
$\mathbb{P}(Z=(j_1,j_2))=u_{j_1,j_2}$.  
(Here $\R_+=[0,\infty), \R_{++}=(0,\infty)$.)
Let $P=(\bp_1,\bp_2)$ and  $\rU(P)$ be the convex set of all probability matrices with marginals $\bp_1,\bp_2$:
\begin{equation}\label{defUPmat}
\rU(P)=\{U=[u_{j_1,j_2}]\in\R_+^{n_1\times n_2}, \sum_{j_2=1}^{n_2} u_{j_1,j_2}=p_{j_1,1}, \quad \sum_{j_1=1}^{n_1} u_{j_1,j_2}=p_{j_2,2}\}.
\end{equation}
Let $C=[c_{i_1,i_2}]\in \R^{n_1\times n_2}$ be the cost matrix of transporting a unit $j_1\in[n_1]$ to $j_2\in[n_2]$.  Then the optimal transport problem is the linear programming problem (LP): 
\begin{equation}\label{TOTmat}
\tau(C,P)=\min\{\langle C,U\rangle, U\in \rU(P)\}. 
\end{equation}
(Here $\langle C,U\rangle=\tr C^\top U$.) 
For $n_1=n_2=n$ and a nonnegative symmetric cost matrix $C=[c_{ij}]$ with zero diagonal satisfying the triangle inequality $c_{ik}\le c_{ij}+c_{jk}$, the quantitty 
 $\tau(C,P)$ gives rise to a distance between probability vectors $\bp_1$ and $\bp_2$, which can be viewed as two histograms. 
It turns out that  $\tau(C,P)$ has many recent applications in machine learning \cite{AWR17,ACB17,LG15,MJ15,SL11}, statistics  \cite{BGKL17,FCCR18,PZ16,SR04} and computer vision \cite{BPPH11,RTG00}.   

A related problem to OT is quantum optimal transport (QOT), see \cite{FECZ22,CEFZ23} and references therein.    QOT is a semidefinite programming problem  \cite{VB96}, which are effectively solved using the interior point methods (ipm) \cite{NN94,Ren01,YFFKNN}.  In this paper we don't treat QOT, but we do use ipm
for solving OT and d-multi-marginal transport problem  that we call $d$-tensor optimal transport abbreviated as $d$-TOT.

Assume that $n_1=n_2=n$.  Then the complexity of finding $\tau(C,P)$ is $O(n^3 \log n)$, as this problem can be stated in terms of flows \cite{PW09}.  In applications, when $n$ exceeds a few hundreds,
the cost is prohibitive.  One way to improve the computation of $\tau(C,P)$ is to replace the linear programming with problem of OT with convex optimization by introducing an entropic regularization term as in \cite{Cut13}.  This regularization terms gives an $\varepsilon$-approximation to $\tau(C,P)$, where $\varepsilon>0$ is given.  The regularization term gives almost linear time approximation $O(n^2)$, ignoring the logarithmic terms, using a variation of the celebrated Sinkhorn algorithm for matrix diagonal scaling \cite{AWR17,LHCJ22,Fri20}.   

The aim of this paper is to introduce the interior point method for $d$-TOT,
 which correspond to the set of $d$-probability measures $\bp_i\in\R^{n_i}$ for $i\in[d]$.   
For the case $d=2$ the set $\rU(P)$ is the set of probability matrices $U\in \R_+^{n_1\times n_2}$ satisfying the marginal conditions $U \1_{n_2}=\bp_1, U^\top \1_{n_1}=\bp_2$.  Here $\1_n\in\R^n$ is the vector whose coordinates are all $1$.
For $d\ge 3$ we introduce $d$-mode tensors $\otimes_{k=1} ^d \R^{n_k}$.  We denote by $\cU\in  \otimes _{k=1}^d \R^{n_k}$ a tensor whose entries are $u_{i_1,\ldots,i_d}$, i.e., $\cU=[u_{i_1,\ldots,i_d}]$.  Assume that $\cC,\cU \in  \otimes _{k=1}^d \R^{n_k}, \cX\in \otimes^{j\in[d]\setminus\{k\}}\R^{n_j}$.   Denote by $\langle \cC,\cU\rangle$ the Hilbert-Schmidt inner product $\sum_{i_k\in[n_k],k\in[d]} c_{i_1,\ldots,i_d}u_{i_1,\ldots,i_d}$.  For $k\in [d]$ denote by  $\y=\cU\times_{\bar k} \cX\in \R^{n_k}$ the contraction on all but the index $k$:
$$y_{i_k}=\sum_{i_j\in[n_j], j\in[d]\setminus\{k\}}u_{i_1,\ldots,i_d} x_{i_1,\ldots,i_{k-1},i_{k+1},\ldots,i_d}.$$
Let $\cJ_{d-1,k}\in\otimes_{j\in [d]\setminus\{k\}}\R^{n_j}$ be the tensor whose all coordinates are $1$.
Define
\begin{equation}\label{defU(P)ten}
\rU(P)=\{\cU\in \otimes_{k=1}^d\R_+^{n_k}, \cU\times_{\bar k} \cJ_{d-1,k}=\bp_k, k\in[d]\},  P=(\bp_1,\ldots,\bp_d).
\end{equation}
Then the tensor optimal transport problem (TOT) is
\begin{eqnarray}\label{TOT}
\tau(\cC,P)\defeq \min\{\langle \cC, \cU\rangle, \cU\in \rU(P)\}.
\end{eqnarray}
TOT problem is a LP problem with $\prod_{k=1}^d n_k$ nonnegative variables and $1+\sum_{k=1}^d (n_k-1)$ constraints. 
The TOT was considered in \cite{Pi68,Po94} in the context of multidimensional assignment problem, where the entires of the tensor $\cU$ are either $0$ or $1$.
There is a vast literature on continuous multidimensional optimal transport problem.
See  for example \cite{FV18,BCN19,TDGU20,HRCK,LHCJ22} and the references therein. 
The TOT problem can be viewed as a discretization the  continuous multidimensional optimal transport problem.

We point out that that the ipm approach is easily adopted for variations of TOT.  Indeed,  for $d>2$ there is another well known variation of the set marginals.  Namely,  let $Z$ be a random variable $Z: \Omega_1\times\cdots \times\Omega_d\to [n_1]\times\cdots\times [n_d]$ with contingency tensor (table)  $\cU\in \R_+^{n_1\times\cdots n_d}$ that gives the distribution of $Z$:
$\mathbb{P}\big(Z=(j_1,\ldots, j_d)=u_{j_1,\ldots,j_d}\big)$.   Denote by $P_k\in \R_+^{n_1\times \cdots n_{k-1}\times n_{k+1}\times \cdots\times n_d}$ the marginal of $Z_k$ obtained from $Z$ with respect to $X_k:\Omega_k\to [n_k]$.    Assume that we are given 
 a distribution of $Z_0$ with a positive contingency tensor $\cV=(v_{j_1,\ldots,j_d})\in \otimes_{i=1}^d\R^{n_i}$ that have the above marginals.  Let $\rU(P)$ be a nonempty set of such $Z$.  See for example \cite{LZZ17} and references therein. Then one can use a similar ipm algorithm to find an approximate algorithm to the problem \eqref{TOT}.
 
 In subsection \ref{subsec:cest} we show that if $\bp_1,\ldots,\bp_d\in\R_+^n$ are weak $K-\ell$ uniform distributions, see Definition \ref{defwKldist}, then the number iterations of the ipm algorithm has complexity $O(n^{d/2})$, ignoring the logarithmic terms.  

\section{The interior point method}\label{sec:ipm}
We first recall some notations and definitions that we will use in this section.
\begin{equation}\label{defellsnrm}
\begin{aligned}
\|\x\|_s=\bigl(\sum_{i=1}^n |x_i|^s\bigr)^{1/s}, \, s\in[1,\infty],\,
\x=(x_1,\ldots,x_n)\in\R^n,\\
\|\x\|\defeq \|\x\|_2,\\
\rB(\x,r)=\{\y\in\R^n, \|\y-\x\|\le r\} \textrm{ for } r\ge 0.
\end{aligned}
\end{equation}

Let $f\in\rC^3(\rB(\x,r))$ for $r>0$.    Denote 
\begin{equation*}
f_{,i_1\ldots i_d}(\x)=\frac{\partial ^d}{\partial x_{i_1}\ldots\partial x_{i_d}}f(\x), \quad i_1,\ldots,i_d\in[n], d\in[3].
\end{equation*}
Recall the Taylor expansion of $f$ at $\x$ of order $3$ for $\bu\in\R^n$ with a small norm:
\begin{equation*}
\begin{aligned}
f(\x+\bu)\approx f(\x)+\nabla f(\x)^\top\bu+\frac{1}{2}\bu^\top \partial^2 f(\x)\bu+\frac{1}{6}\partial ^3 f(\x)\otimes \bu^{3\otimes},\\
\nabla f(\x)=(f_{,1}(\x),\ldots,f_{,n}(\x))^\top,  \quad \partial^2 f(\x)=[f_{,ij}(\x)],  i,j\in[n],\\ \partial^3 f(\x)=[f_{,ijk}(\x)], i,j,k\in[n], \, \partial^3 f(\x)\otimes\bu^{3\otimes}=\sum_{i,j,k\in[n]}f_{,ijk}(\x)u_i u_j u_k, 
\end{aligned}
\end{equation*}
where $\nabla f, \partial^2 f, \partial ^3 f$ are called
the gradient, the Hessian, and the 3-mode symmetric partial derivative tensor of $f$. 
A set $\rD\subset \R^n$ is called a domain if $\rD$ is an open connected set.
\begin{definition}\label{defconcconst}  
Assume that $f: \rD\to\R$ is a convex function 
 in a convex domain $\rD\subset \R^n$,  and $f\in\rC^3(\rD)$.   The function $f$ is called   $a (>0)$-self-concordant, or simply self-concordant, if the following inequality hold
\begin{equation}\label{defconcconst1} 
|\langle \partial^3f(\x),\otimes^3\bu\rangle|\le 2a^{-1/2} (\bu^\top \partial^2f(\bx)\bu)^{3/2}, 
\textrm{ for all }\x\in \rD,\bu\in\R^n.
\end{equation}
The function $f$ is called a standard self-concordant if $a=1$,  and a strongly $a$-self-concordant if $f(\x_m)\to\infty$ if the sequence $\{\x_m\}$ converges to the boundary of $\rD$.

The complexity value $\theta(f)\in[0,\infty]$ of an a-self-concordant function $f$ in $\rD$, called a self-concordant parameter in \cite[Definition 2.3.1]{NN94}, is
\begin{equation}\label{defsconc0}
\begin{aligned}
\theta(f)=sup_{\x\in\rD}
\inf\{\lambda^2\in[0,\infty], |\nabla f(\x)^\top \bu|^2\\
\le \lambda^2 a\big(\bu^\top\partial^2 f(\x)\bu\big),\forall \bu\in\R^n\}.
\end{aligned}
\end{equation}
 A strongly self-concordant function with a finite $\theta(f)$ is called a barrier (function).
\end{definition}

The following lemma is probably well known, and we give its short proof for completeness:
\begin{lemma}\label{scfstr}  Let $\rD\subset\R^n$ be a convex domain and assume that $f$ is a self-concordant function in $\rD$.  Then one of the following conditions hold
\begin{enumerate}[(a)]
\item The function $f$ is affine on $\rD$.
\item The Hessian $\partial^2 f$ is positive definite on $\rD$.
\item There is an orthogonal change of coordinates $\y=(y_1,\ldots,y_n)^\top=Q\x$, such that 
\begin{equation*}
f(\y)=f_1((y_1,\ldots,y_m)^\top)+by_{m+1}
\end{equation*} 
for some $m\in[n-1]$, such that $f_1$ has a positive definite Hessian on $\rD_1\subset \R^m$, where $\rD_1$ is the projection of $\rD$ on the first $m$ coordinates.
\end{enumerate}
Suppose furthermore that $\theta(f)<\infty$.   Then either $f$ is constant on $\rD$,  the Hessian of $f$ is positive definite in $\rD$, or the condition (c) holds with $b=0$.
In particular, $\nabla f(\x)$ is orthogonal to the kernel
of $\partial^2 f(\x)$ for $\x\in\rD$.
\end{lemma}
\begin{proof}
Corollary 2.1.1 in \cite{NN94} states that the nullity subspace $\U(\x)\subseteq\R^n$ of $\partial^2 f(\x),\x\in\rD$ does not depend on $\x\in\rD$.  Set $\U=\U(\x)$ for $\x\in\rD$.  If $\U=\R^n$ then (a) holds.   If $\U=\{\0\}$ then (b) holds.  If $\dim  \U=n-m$ it is straightforward to show that (c) holds. 

Assume that $\theta(f)<\infty$.   Suppose that $\0\ne\bu\in \ker \partial^2 f(\x)$.  Then $\nabla f(\x)^\top \bu=0$.   If the condition (a) satisfied then $f$ is a constant function.  Suppose that the condition (c) is satisfied.  Then $b=0$.
In particular, $\nabla f(\x)$ is orthogonal to the kernel
of $\partial^2 f(\x)$ for $\x\in\rD$.
\end{proof}

Denote by $\rS_n\supset\rS_{n,+}\supset \rS_{n,++}$ the space of $n\times n$ symmetric, the cone of positive semidefinite and the open set of positive definite matrices respecctively.   For $A,B\in \rS_n$ we denote $A\succ B\,(A\succeq B)$ if $A-B\in\rS_{n,++},\, (A-B\in\rS_{n,+})$.
For a matrix $A\in\R^{n\times n}$ denote by  $A^\dagger\in\R^{n\times n}$ the Moore-Penrose inverse of $A$ \cite[\textsection 4.12]{Frib}. Recall that if $A$ is invertible then $A^\dagger=A^{-1}$.  In particular for $a\in\R$: $a^\dagger =a^{-1}$ if $a\ne 0$, and $a^\dagger =0$ if $a=0$.
Assume that $A\in \rS_n$.  Then $A=Q\diag(\lambda_1,\ldots,\lambda_n) Q^\top$, where $Q\in\R^{n\times n}$ is an orthogonal matrix and
\begin{equation*}
\lambda_{\max}=\lambda_1\ge\ldots\ge\lambda_n=\lambda_{\min}
\end{equation*} 
are the eigenvalues of $A$. Then $A^\dagger=Q \diag(\lambda_1^\dagger,\ldots,\lambda_n^\dagger) Q^\top$.   In particular, $\ker A=\ker A^\dagger$.  In what follows we will use the following lemma:
\begin{lemma}\label{maxcharlem}
Let $A\in\rS_{n,+}, \y\in\R^{n}$.  Suppose that $\y^\top \ker A=0$.  Then
\begin{equation}\label{maxcharlem1}
\begin{aligned}
\y^\top A^\dagger\y=\inf\{\lambda>0, |\y^\top\bu|^2\le \lambda^2 \bu^\top A\bu, \\
\forall \bu\in\R^n\}= \max_{\bu\in\R^n} 2\y^\top \bu -\bu^\top A\bu.
\end{aligned}
\end{equation}
Furthermore,  if $A\succeq \y\y^\top$ then 
$\y^\top A^\dagger\y \le 1$.
\end{lemma}
\begin{proof} Clearly, it is enough to consider the case $\y\ne \0$. 
Let 
$$\mu=\inf\{\lambda>0, |\y^\top\bu|^2\le \lambda^2 \bu^\top A\bu, \forall \bu\in\R^n\}.$$
Suppose first that $A$ is positive definite.  Then  $\mu=\max_{\bu\ne \0}\big(\bu^\top\by \by^\top \bu\big)/\big(\bu^\top A\bu\big)$.  Let $B=\sqrt A\succ 0$, and $ B\bu=\bv$.   Then $\mu$ is the maximum eigenvalue of the rank-one matrix $B^{-1} \y\y^\top B^{-1}$.
Thus 
$$\mu=\tr B^{-1} \y\y^\top B^{-1}=\y^\top B^{2}\y=\y^\top A^{-1}\y.$$ 
This proves the first equality in \eqref{maxcharlem}.  

We now show the second equality in \eqref{maxcharlem}.  
Fix $\bw\ne \0$ and let $\bu=t\bw$.  Set  $\phi(t)=2\y^\top(t\bw)-(t\bw)^\top  A(t\bw)$.  The  maximum of $\phi(t)$ is achieved at $t=\frac{\y^\top\bw}{\bw^\top A\bw}$ and is equal to $\frac{|\y^\top \bw|^2}{\bw^\top A\bw}$.   Use the first equality of \eqref{maxcharlem} to deduce the second equality in \eqref{maxcharlem1}.

Assume now that $A\in\rS_{n,+}$ is singular.
Suppose that $\bu^\top A\bu=0$.  Then $\bu\in\ker A$.  Hence, $\y^\top \bu=0$.  Therefore, its is enough to consider the case where $\bu\in$range$\, A$.   As $\y^\top\ker A=0$ it follows that $\y\in$range$\,A$.  Let $C$ be the restriction of $A$, viewed as a linear operator $\R^n\to\R^n$,  to range~$A$.  So $C$ is positive definite and we can use the previous case.
Observe that $\y^\top C^{-1}\y=\y^\top A^\dagger\y$, and the first equality of  \eqref{maxcharlem} follows.  The second equality follows similarly.  

Suppose that $A\succeq \y\y^\top$.  That is, $A=\y\y^\top +B$ for some $B\in\rS_{n,+}$.
Hence $\ker A=\ker (\y\y^\top)\cap\ker B$.  Therefore $\y$ is orthogonal to $\ker A$.
We now use the second equality in \eqref{maxcharlem1}.  Observe that $\bu^\top A\bu\ge \bu^\top (\y\y^\top)\bu$.   Hence
\begin{equation*}
\y^\top A^\dagger \y\le \y^\top (\y\y^\top)^\dagger\y=1.
\end{equation*}
\end{proof}

\begin{corollary}\label{charthetf}  Let $\rD\subset\R^n$ be a convex domain and assume that $f$ is a nonconstant $a$-self-concordant function in $\rD$.    Suppose furthermore that $\nabla f(\x)^\top \ker \partial^2 f(\x)=0$ for each $\x\in\rD$.   Then
\begin{equation}\label{defconcconst1}
 \theta(f)=a^{-1}\sup_{\x\in \rD}\nabla f(\x)^\top (\partial^2f)^\dagger(\x)\nabla f(\x).
\end{equation}
\end{corollary}
This equality is well known if $\partial^2f$ is invertible \cite[top of page 16]{NN94}.
A simple example of a strongly standard self-concordant function for the interior of the cone $\R_+^n$,  denoted as $\R_{++}^n$, where $\R_{++}=(0,\infty)$, is 
\begin{equation}\label{logxbar}
\begin{aligned}
\sigma(\x)=-\sum_{i=1}^n \log x_i,\\
\theta(\sigma)=n.
\end{aligned}
\end{equation}
The equality $a=1$ follows from the well known fact that the norm $\|\x\|_s$ is decreasing for $s\in[1,\infty]$.  Use Corollary \ref{charthetf} 
to deduce the second equality of \eqref{logxbar}.

 Assume that $\rD$ is a bounded convex domain,  $\x\in\rD$  and $\rL$ is a line  through $\x$. 
 Denote by $d_{\max}(\x,\rL)\ge d_{min}(\x,\rL)$ the two distances from $\x$ to the end points of $\rL\cap\partial\rD$.   Then $\textrm{sym}(\x,\rD)$ is the infimum of $\frac{d_{\min}(\x,L)}{ d_{\max}(\x,L)}$ for all lines $\rL$ through $\x$.   Observe that if $\rB(\x,r)\subset \textrm{Closure}(\rD)\subset \rB(\x,R)$ then $\textrm{sym}(\x,\rD)\ge r/R$.

Recall that Renegar \cite{Ren01} deals only with strongly standard self-concordant functions.  
The complexity value $\theta(f)$, coined in \cite{Ren01}, is called the parameter of barrier $f$ in \cite{NN94}, and is considered only for self-concordant barrier  in \cite[\textsection 2.3.1] {NN94}.   

We now recall the  complexity result to approximate the infimum of a linear functional on a bounded convex domain with whose boundary is given by a barrier function $\beta$.
We normalize $\beta$ by assuming that it is strongly self-concordant.
A simple implementation of the Newton's method  is  a "short-step" ipm's that  follows the central path \cite[\textsection 2.4.2]{Ren01}.
The number of iterations to approximate the minimum of a linear functional  within $\varepsilon$ precision starting with an intial point $\x'$ is \cite[Theorem 2.4.1]{Ren01}:
\begin{equation}\label{Renthm}
O\big(\sqrt{\theta(\beta)}\log\big(\frac{\theta(\beta)}{\varepsilon \textrm{sym}(\x',\rD)}\big)\big).
\end{equation}
\subsection{The number of iterations of ipm for matrix optimal transport}\label{subsec:mot}
Assume that $1<m,n\in\N$.
Let $\bp=(p_1,\ldots,p_m)^\top,\bq=(q_1,\ldots,q_n)^\top$ be two positive probability vectors.  Denote by $\rU(P)$ the set \eqref{defUPmat}, where $n_1=m, n_2=n$ and $\bp_1=\bp, \bp_2=\bq$.   Set $\1_m=(1,\ldots,1)^\top\in\R^m$, and define
\begin{equation*}
\rU_{0,2}=\{X=[x_{ij}]\in\R^{m\times n},  X\1_n=\0, X^\top \1_m=\0\}.
\end{equation*}
As the sum of all rows of $X$ is equal to the sum of all columns of $X$ it follows that $\dim\rU_{0,2}=(m-1)(n-1)$.  
Let $\be_i=(\delta_{1i},\ldots,\delta_{mi})^\top\in\R^m, i\in[m]$ and $\bbf_j=(\delta_{1j},\ldots,\delta_{nj})^\top\in\R^n, j\in[n]$ be the standard bases in $\R^m$ and $\R^n$ respectively.  Then one has a following simple basis in $\rU_{0,2}$:
\begin{equation*}
\begin{aligned}
\bg_i\bh_j^\top, i\in[m-1], j\in[n-1],\\
\bg_i=\be_i-\be_{i+1}, i\in[m-1],\quad \bh_j=\bbf_j-\bbf_{j+1},j\in[n-1].
\end{aligned}
\end{equation*}

The interior of $\rU(P)$, denoted as $\rU_o(P)$,  is the set of positive matricers in the affine space 
$$\rU_{af,2}(P)=\{X=\bp\bq^\top +\sum_{i=1}^{m-1}\sum_{j=1}^{n-1} t_{ij}\bg_i\bh_j^\top, \quad t_{ij}\in\R,  i\in[m-1],j\in[n-1]\}. $$
Let  
$$\sigma(X)=-\sum_{i=1}^m\sum_{j=1}^n \log x_{ij}, \quad X=[x_{ij}]\in\R_{++}^{m\times n}$$
be a barrier function $\R_{++}^{m\times n}$.  Recall that $\sigma$ is a standard self-concondant barrier with $\theta(\sigma)=mn$.   The restriction of $\tilde \sigma$ to $\rU_{af,2}\cap \R_{++}^{m\times n}$ is a standard self-concordant barrier with 
\begin{equation}\label{thetsig}
\theta(\tilde\sigma)\le mn.
\end{equation}
\begin{theorem}\label{ipmotmat}
Let $\bp=(p_1,\ldots,p_m)^\top\in\R^m,\bq=(q_1,\ldots,q_n)^\top\in\R^n$ be positive probability vectors.  Consider the minimum problem \eqref{TOTmat} on the polytope
 $\rU(P)$ given by \eqref{defUPmat}, where $n_1=m, n_2=n$ and $\bp_1=\bp, \bp_2=\bq$. 
The short step interior path algorithm with the barrier $\tilde \sigma$ starting at the point $\bp\bq^\top$ finds the value $\tau(C,P)$ within precision $\varepsilon>0$ in  
\begin{equation}\label{ipmotmat1}
O\big(\sqrt{mn}\log\frac{\sqrt{2}mn}{\varepsilon (\min_{i\in[m]}p_i)(\min_{j\in[n]}q_j)}\big)
\end{equation}
iterations.
\end{theorem}
\begin{proof} In view of \eqref{Renthm} and \eqref{thetsig}  it is enough to show that 
$$\textrm{sym}(\bp\bq^\top,\rD)\ge (\min_{i\in[m]}p_i)(\min_{j\in[n]}q_j)\big)/\sqrt{2},\quad
\rD=\rU_{af}(P)\cap \R_{++}^{m\times n}.$$ 
For $Y=[y_{ij}]\in \R^{m\times n}$ denote: $\|Y\|_2=\sqrt{\sum_{i=1}^m\sum_{j=1}^n y_{ij}^2}$.
Assume that $X\in \textrm{Closure}(\rD)$.   Then  $X=[x_{ij}]$ is a  probability matrix.  
Note that if $X\in\partial \rD$ then $x_{ij}=0$ for some $i\in[m],j\in[n]$. Hence,  for $X\in\partial \rD$ we have the inequalities:
\begin{equation*}
\begin{aligned}
\sqrt{2}\ge\|\bp\bq^\top-X\|_2\ge p_iq_j   \textrm{ for some } i\in[m],j\in[n]\Rightarrow\\ 
 \sqrt{2}\ge d_{\max}(\bp\bq^\top,L)\ge d_{\min}(\bp\bq^\top,L)\ge(\min_{i\in[m]}p_i)(\min_{j\in[n]}q_j)\big).
\end{aligned}
\end{equation*}
\end{proof}
\subsection{The number iterations of ipm for tensor optimal transport}\label{subsec:totipm}
Assume that $d>2$.
We first consider the TOT of the form \eqref{TOT}.  We now repeat the arguments of the previous subsection.  Let 
\begin{equation*}
\begin{aligned}
\rU_{0,d}=\{\cU\in \otimes^d\R^n, \cU\times_{\bar k} \cJ_{d-1}=0, k\in[d]\},,\\
\rU_{af,d}(P)=\otimes_{k=1}^d \bp_k +\rU_{0,d}, \quad P=(\bp_1,\ldots,\bp_d)
\end{aligned}
\end{equation*}
Then $\rU(P)=\rU_{af,d}(P)\cap \otimes_{k=1}^d \R_+^{n_k}$.  The interior of $\rU(P)$ is given by $\rU_{af,d}\cap \otimes_{k=1}^d \R_{++}^{n_k}$.  Let
$$\sigma(\cX)=-\sum_{i_k\in[n_k], k\in[d]}\ \log x_{i_1,\ldots,i_d}, \quad \cX\in\otimes_{k=1}^d\R_{++}^{n_k}$$
be a barrier function on $\otimes_{k=1}^d\R_{++}^{n_k}$.   Thus, $\sigma$ is a standard self-concondant barrier with $\theta(\sigma)=\prod_{k=1}^d n_k$.   The restriction of $\tilde \sigma$ to $\rU_{af,2}\cap \R_{++}^{m\times n}$ is a standard self-concordant barrier with 
\begin{equation}\label{thetsigd}
\theta(\tilde\sigma)\le \prod_{k=1}^d n_k.
\end{equation}
The arguments of the proof of Theorem \ref{ipmotmat} yield:
\begin{theorem}\label{ipmotd}
Let $\bp_k=(p_{1,k},\ldots,p_{n_k,k})^\top\in\R^{n_k}, k\in[d]$ be positive probability vectors.  Consider the minimum problem \eqref{TOT} on the polytope
 $\rU(P)$ given by \eqref{defU(P)ten}. 
The short step interior path algorithm with the barrier $\tilde \sigma$ starting at the point $\otimes_{k=1}^d\bp_k$ finds the value $\tau(C,P)$ within precision $\varepsilon>0$ in  
\begin{equation}\label{ipmotmat1}
O\big(\sqrt{\prod_{k=1}^d n_k}\log\frac{\sqrt{2}\prod_{k=1}^d n_k}{\varepsilon \prod_{k=1}^d\min_{i_k\in[n_k]}p_{i_k,k}}\big)
\end{equation}
iterations.
\end{theorem}

We now consider a variation of the polytope $\rU(P)$, which correspond to the problem of $d$-dimensional stochastic tensors \cite{LZZ17}.   For $\cU=[u_{i_1,\ldots,i_d}]\in \otimes_{k=1}^d \R^{n_k}$ and a vector $\x=(x_1,\ldots,x_{n_k})^\top\in\R^{n_k}$ denote 
\begin{equation*}
\begin{aligned}
\cU\times _k \x =\cW=[w_{i_1,\ldots,i_{k-1},i_{k+1},\ldots,i_d} ]\in \otimes_{j\in[d]\setminus\{k\}}\R^{n_j},\\
w_{i_1,\ldots,i_{k-1},i_{k+1},\ldots,i_d}=\sum_{i_k=1}^{n_k} u_{i_1,\ldots,i_d} x_{i_k}.
\end{aligned}
\end{equation*}
Define
\begin{equation}\label{defV(P)ten}
\begin{aligned}
\rV(P)=\{\cV=[v_{i_1,\ldots,i_d}]\in \otimes_{k=1}^d\R_+^{n_k}, \\
\cV\times_{k} \1_{n_k}=\otimes_{j\in[d]\setminus{k}}\bp_j, k\in[d]\},  P=(\bp_1,\ldots,\bp_d),\\
\end{aligned}
\end{equation}
Let 
\begin{equation}\label{defV0(P)}
\begin{aligned}
\rV_{0,d}=\{\cV=[v_{i_1,\ldots,i_d}]\in \otimes_{k=1}^d\R^{n_k}, 
\cV\times_{k} \1_{n_k}=0\in[d]\},\\
\rV_{af,d}(P)=\otimes_{k=1}^d \bp_k+\rV_{0,d}.
\end{aligned}
\end{equation}
Let $\be_{1,k},\ldots,\be_{n_k,k}\in\R^{n_k}$ be the standard basis in $\R^{n_k}$ for $k\in[d]$.   Denote
\begin{equation*}
\bg_{i,k}=\be_{i,k}-\be_{i+1,k} \textrm{ for } i\in[n_k-1].
\end{equation*}
Observe that $\1_{n_k}^\perp=$span$(\bg_{1,k},\ldots,\bg_{n_k-1,k})$ is the orthogonal
complement of $\1_{n_k}$ in $\R^{n_k}$.
We claim that 
\begin{equation}\label{V0dstr}
\rV_{0,d}=\otimes_{k=1}^d\1_{n_k}^\perp.
\end{equation}
Indeed, assume that $\cV\in\otimes_{k=1}^d \R^{n_k}$ satisfies $\cV\times_1\1_{n_1}=0$.  View $\cV$ as a matrix in $\R^{n_1}\otimes\big(\otimes_{k=2}^d\R^{n_k}\big)$.  The above condition  yields that range~$ \cV\subset \1_{n_1}^\perp$, which is equivalent to $\cV\in \1_{n_1}^\perp\otimes\big(\otimes_{k=2}^d \R^{n_k}\big)$.  Apply this observation to $\rV_{0,d}$ to deduce \eqref{V0dstr}.  Hence,
\begin{equation*}
\rV_{af,d}(P)=\{\cV=\otimes_{k=1}^d \bp_k +\sum_{i_k\in[n_k-1],k\in[d]} t_{i_1,\ldots,i_d}\otimes_{j=1}^d \bg_{i_j,j}, [t_{i_1,\ldots,i_d}]\in\otimes_{l=1}^d \R^{n_l-1}\}. 
\end{equation*}
Then $\rV(P)=\rV_{af,d}(P)\cap \otimes_{k=1}^d \R_+^{n_k}$.  The interior of $\rV(P)$ is given by $\rV_{af,d}\cap \otimes_{k=1}^d \R_{++}^{n_k}$.  

The arguments of the proof of Theorem \ref{ipmotmat} yield:
\begin{theorem}\label{ipmotdv}
Let $\bp_k=(p_{1,k},\ldots,p_{n_k,k})^\top\in\R^{n_k}, k\in[d]$ be positive probability vectors.  Consider the minimum problem $
\tau(\cC,P)= \min\{\langle \cC, \cU\rangle, \cU\in \rV(P)\}$.
The short step interior path algorithm with the barrier $\tilde \sigma$ starting at the point $\otimes_{k=1}^d\bp_k$ finds the value $\tau(C,P)$ within precision $\varepsilon>0$ in the number of iterations given by \eqref{ipmotmat1}.
\end{theorem}
\subsection{Iteration estimates for certain probabilities $P$}\label{subsec:cest}
Our iteration estimate \eqref{ipmotmat1} depends on $P=(\bp_1,\ldots,\bp_d)$: the product of the minum values of the coordinates of $\bp_k$ for $k\in[d]$.   
\begin{definition}\label{defwKldist}
A probability vector $\bp=(p_1,\ldots,p_n)\in\R_{+}^n$ is called a weak $K-\ell$ uniform distribution if
\begin{equation}\label{defwKldist1}
p_j\ge \frac{K}{n^{\ell}}, j\in[n],\quad K>0, \ell\ge 1, K\le n^{\ell-1}.
\end{equation}
\end{definition}
Note that if $K=\ell=1$ the $\bp$ is the uniform distribution.

Theorem \ref{ipmotd} yields:
\begin{corollary}\label{cipmotd}  Let the assumptions Theorem \ref{ipmotd} hold.
Assume that each $\bp_j$ is a weak $K-\ell$ uniform distribution.  Then the short step interior path algorithm with the barrier $\tilde \sigma$ starting at the point $\otimes_{k=1}^d\bp_k$ finds the value $\tau(C,P)$ within precision $\varepsilon>0$ in  
\begin{equation}\label{cipmotmat1}
O\big(\sqrt{\prod_{k=1}^d n_k}\log (\sqrt{2}\varepsilon^{-1}K^{-d}\prod_{k=1}^d n_k^{1+\ell})\big)
\end{equation}
iterations. In particular, if $n_1=\cdots=n_d=n$ then the above estimate is $O\big(n^{d/2}\log(\sqrt{2}\varepsilon^{-1} K^{-d} n^{d(1+\ell)})\big)$. 
\end{corollary}
 \section*{Acknowledgment}
The author is partially supported by the Simons Collaboration Grant for Mathematicians.  
\bibliographystyle{plain}

\end{document}